OpenStudy (anonymous):

What is 247 grams converted to kilograms?

jimthompson5910 (jim_thompson5910):

1000 grams = 1 kilogram 1000/1000 grams = 1/1000 kilogram 1 gram = 0.001kilogram

jimthompson5910 (jim_thompson5910):

Now multiply both sides by 247

OpenStudy (anonymous):

247000 grams?

jimthompson5910 (jim_thompson5910):

1 gram = 0.001kilogram 247*1 gram = 247*0.001kilogram

jimthompson5910 (jim_thompson5910):

247 times 1 = 247, so the left side is 247 grams 247 times 0.001 = ???

OpenStudy (anonymous):

.247 ?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

Thanks jim

jimthompson5910 (jim_thompson5910):

yw